### v1

# DrishtiKon: Multi-Granular Visual Grounding for Text-Rich Document Images

Abstract
Visual grounding in text-rich document images is a critical yet underexplored challenge for document intelligence and visual question answering (VQA) systems. We present DhrishtiKon, a multi-granular visual grounding framework designed to enhance interpretability and trust in VQA for complex, multilingual documents. Our approach integrates robust multi-lingual OCR, large language models, and a novel region matching algorithm to accurately localize answer spans at block, line, word, and point levels. We curate a new benchmark from the CircularsVQA test set, providing fine-grained, human-verified annotations across multiple granularities. Extensive experiments demonstrate that our method achieves state-of-the-art grounding accuracy, with line-level granularity offering the best trade-off between precision and recall. Ablation studies further highlight the benefits of multi-block and multi-line reasoning. Comparative evaluations with leading vision-language models reveal the limitations of current VLMs in precise localization, underscoring the effectiveness of our structured, alignment-based approach. Our findings pave the way for more robust and interpretable document understanding systems in real-world, text-centric scenarios. Code and dataset has been made available at https://github.com/kasuba-badri-vishal/DhrishtiKon.
1 Introduction
Visual Grounding is the task of locating specific visual elements in an image that correspond to a given text query/question in natural language as input to a corresponding probable answer that answers the query from the input image[8]. This task is particularly relevant in the context of Visual Question Answering (VQA) task, where the goal is to provide accurate and contextually relevant answers based on the visual content of an image, providing both transparency and trust on top of the predicted answer’s relavance within the image. Also, Grounded learning for the task allows for more accurate and contextually relevant responses, as it requires models to focus on the visual content directly related to the question to finding the grounded answers. In this work, we will explore Visual Grounding task on text-centric images and also evaluate the performance on a constructed multi-granular grounding benchmark from the collection of Government Circular images. We propose a Multi-granular Visual Grounding framework, DhrishtiKon which aims to enhance the model’s ability to ground answers in visual elements at multiple levels of granularity, supporting multiple pages which utilized a novel region matching algorithm to improve the grounding process.
2 Literature Study
The rise of Multimodal Large Language Models (MLLMs) has advanced document understanding, yet visual grounding in text-rich document images remains an underexplored challenge but yet vital for document intelligence tasks. Recent efforts such as TGDoc[7], DOGE-Bench[9], and TRIG-Bench[3] have introduced benchmarks and models to address this gap. The Table 1 compares these benchmarks across data, methodologies, support and synthetic data generation strategies.
Table 1: Comparison of Visual Grounding Benchmarks on Text-Rich Images
Benchmarks Image Source Quantity Granularity Support Synthetic data generation OCR LLM TGDoc Online Powerpoint samples 99k Slides 12k high quality Chats Single Single-Block PaddleOCR GPT4 DOGE-Bench Charts, Posters, PDFs 700k Instruction 4k Benchmark Multi-Granular Single-Block EasyOCR GPT4o TRIG-Bench DocVQA, InfographicVQA 90k Synthetic 800 Manual Single Multi-Block PaddleOCR GPT4o
TGDoc[7] improves spatial awareness in MLLMs through instruction tuning on 99K PowerPoint slides and 12K GPT-4-generated dialogues, using BLIP2, PaddleOCR, and GPT-4 to align text recognition with spatial cues. DOGE-Bench[9] introduces multi-granular grounding and referring tasks across charts, posters, and PDFs, using a synthetic pipeline (DOGE-Engine) to produce 700K training samples and a 4K evaluation benchmark, enabling fine-grained referencing at block, line, and word levels. In contrast, TRIG-Bench[3] emphasizes grounding in real-world, layout-heavy documents using a hybrid OCR-LLM-human pipeline, offering 90K synthetic and 800 manually curated samples to assess MLLM performance in spatially complex settings. Collectively, these benchmarks expose the limitations of current MLLMs in handling dense layouts and highlight the need for improved spatial reasoning in document-centric tasks. While these benchmarks have made significant strides in visual grounding, they also reveal critical gaps in the current state of MLLMs. The key gaps remain particularly in multilingual, multi-block scenarios with human-verified annotations applicable to all types of text-rich images. In this thesis, we aim to develop multilingual, multi-block and multi-granular visual grounding benchmark that address these limitations and could pave the way for more robust and versatile document understanding systems.
3 DhrishtiKon Methodology
This section details the proposed pipeline for DhrishtiKon, a multi-granular visual grounding solution for text-rich document images. Our approach integrates multi-lingual OCR[4, 5], large language models (LLMs) [2], and a hybrid scoring-based region matching algorithm as discussed in Section 3.4 for accurate localization of semantically relevant text blocks. Figure 1 provides an overview of the DhrishtiKon pipeline.
Figure 1: DhrishtiKon: Multi-granular Visual Grounding Solution Pipeline with matching algorithm flowchart on an example Illustration VQA sample.
3.1 Input Acquisition
We begin with an input image of a text-rich document, typically a scanned official circular or memorandum, which forms the visual basis for question answering. Accompanying the image is a natural language question that queries for any extractive information, such as specific data, references, or contextual details present in the document.The question is designed to be open-ended, allowing for a range of possible answers that may span multiple text blocks within the document, but extractive in nature intending to find specific phrases or data points within the image content.
3.2 Multi-Lingual Block-Level OCR
The input image is processed using a robust multi-lingual OCR system using DocTR [4] and Surya-OCR[5] that segments the document into block-level text regions via Layout prediction. Each region is encoded with its corresponding bounding box coordinates and transcribed content. This enables spatial decomposition of the document for fine-grained semantic grounding.
3.3 Question-Conditioned Answer Prediction
The extracted blocks, along with the user-input question, are fed into a large language model (LLM). We use LlaMA-3.1-8B instruct open-source model[2] as our model, which performs semantic reasoning over the textual content and outputs a predicted answer. This step simulates an MLLM (Multimodal Large Language Model) inference in purely textual modality, bypassing the need for visual input during answer generation as it takes OCR output as contextual information.
3.4 Candidate Region Matching Algorithm
To locate the predicted answer within the document, we implement a matching algorithm that loops over each probable text block extracted by the OCR engine. For each block, a composite matching score is calculated using the following components:
  1. 1.
    Fuzzy Score: Computed using both the partial match and token-based matching similarity between the predicted answer and block content.
  2. 2.
    Length Factor: Rewards blocks with reasonable length relative to the predicted answer with comparison of probable block matched text.
  3. 3.
    Penalty Functions: Down-ranks blocks with either excessively short bounding boxes (indicative of noise) or lacking contextual semantic overlap.
As described in the Algorithm 1, Each candidate region is scored based on components which inturn have a scale of contribution, and those regions that scored above a pre-defined threshold are retained. The candidate blocks are sorted based on their composite scores. The top-k matched blocks are selected and visualized along with the predicted answer for interpretability. This allows the luxury to choose the number of multi-block predictions required based on the threshold parameters.
Algorithm 1 Region Matching Algorithm for Block-Level Probable Regions
1:Input:
2:   1. Question Text $Q$
3:   2. Predicted Answer Text $A$
4:   3. Block-level regions $R=\{r_{1},r_{2},\dots,r_{n}\}$
5:for each region $r\in R$ do
6:     Step 1: Compute Fuzzy Score
7:        a. Via partial match between $A$ and text in $r$
8:        b. Via token match between $A$ and text in $r$
9:     Step 2: Compute Length Factor
10:        Compute ratio or penalty based on bounding box/text length
11:     Step 3: Compute Answer Match Score
12:        Evaluate similarity between $A$ and content of $r$
13:     Step 4: Apply Penalties
14:        a. Penalize for short bounding boxes
15:        b. Penalize for non-contextual or irrelevant text
16:     Step 5: Compute Combined Score
17:     if Combined Score $\geq$ Threshold then
18:         Mark $r$ as valid match
19:     end if
20:end for
21:Step 6: Sort matched regions based on Combined Score
22:Step 7: Return top $K$ matches
4 Experiments and Results
4.1 Grounding Datasets
To evaluate the effectiveness of our visual grounding framework, we curate a benchmark dataset derived from the Indic Government Circulars. We sample 70 diverse document images encompassing complex layout structures and multi-lingual textual content, making them suitable for evaluating fine-grained spatial-semantic reasoning. The dataset includes detailed annotations as summarized in Table 2. As the TGDoc[7], DOGE-Bench[9] and TRIG-Bench[3] datasets which are the existing Visual Grounding on text-centric images are yet not released and available as described in Section 2 , we perform our experiments testing on the curated testset comprising of 509 QnA pairs with grounding.
Table 2: Statistics of the annotated dataset from CircularsVQA testset for Multi-Granular Visual Grounding.
Attribute Count
Document Images 70
Question–Answer (QnA) Pairs 509
Block-level regions 538
Line-level regions 988
Word-level regions 5,968
Point-level data 538
4.1.1 Annotation Tool
Figure 2: In-house annotation tool supporting multi-level region labeling.
To construct the annotated test set used in our experiments, we developed a custom in-house annotation tool designed to support hierarchical and fine-grained annotation of document regions. This tool enables annotation at the block, line, word, and point levels, offering flexibility in defining answer spans that match natural visual layouts. The tool has an interactive interface (Figure 2) that allowed to label precise spans and relationships and also supports automatic text detection at word-level from the line-level annotated region aiding in enhanced annotation. Unlike other existing datasets[7, 9], the dataset has been manually annotated instead of using synthetic data generation straties via properietary LLM.
Figure 3: Visual grounding results across different granularities: Block, Line, Word, and Point. The bounding boxes indicate the predicted regions corresponding to the question and answer.
4.2 Granularity-wise Evaluation
We evaluate the performance of our proposed visual grounding pipeline using standard metrics: Precision, Recall, and F1-score. These metrics are computed across four levels of grounding granularity: Block, Line, Word, and Point. The results, presented in Table 3, correspond to the predictions generated using our Region Matching Algorithm under the Question + Ground Truth Answer input. Figure 3 illustrates the visual grounding qualititative results across these granularities.
Table 3: Evaluation of grounding accuracy at different granularities
Level Precision (%) Recall (%) F1-score (%)
Block 53.89 68.22 60.21
Line 65.06 73.68 69.10
Word 60.32 49.87 54.60
Point 60.66 51.30 55.79
We observe that Line-level grounding achieves the highest F1-score of 69.10%, demonstrating that this granularity provides the best trade-off between precision and recall. This is likely due to the structural nature of government circulars, where answers often span single, well-formed lines. Block-level grounding benefits from capturing semantically complete units but suffers from region fragmentation and inconsistent bounding box aggregation. Meanwhile, word-level and point-level grounding, although more fine-grained, are more susceptible to errors from OCR misalignment, which leads to reduced recall. Overall, these findings suggest that mid-level granularities (especially lines) are optimal for visual grounding in structured documents. However, as highlighted in recent work such as MOLMO[1], point-level finetuning can still be beneficial for training as it provides required minimal grounding context and reducing training costs.
4.3 Model-wise Evaluation
We further compare the Region Matching Algorithm with multiple vision-language models (VLMs), including LLaMA 3.1[2] and Qwen2.5-VL[6]. All models are evaluated on the line-level grounding task under various supervision regimes: using both the question and predicted answer (Q+A), using only the question, and hybrid configurations combining predicted answers with grounding algorithms. The results are presented in Table 4.
Table 4: Line-level grounding evaluation (10 max boxes, IoU = 0.5) with OCR-based input (text + bbox).
Input (Text + BBox) OCR (QA) Model (QA) OCR (Grnd.) Model (Grnd.) P (%) R (%) F (%)
Ground Truth Answer - - YES Algorithm 65.06 73.68 69.10
Ground Truth Answer - - YES LLAMA 49.20 64.98 56.00
Ground Truth Answer - - No Qwen2.5VL 6.00 4.55 5.18
Predicted Answer YES LLaMA YES Algorithm 43.97 53.14 48.12
Predicted Answer YES LLaMA YES LLAMA 37.43 47.77 41.97
Predicted Answer NO PATRAM YES Algorithm 35.58 30.97 33.12
The Region Matching Algorithm, when paired with ground truth answers, clearly outperforms all other methods and could be considered as the skyline performance for line-level grounding tasks in this dataset as the ideal conditions are met.
Qwen2.5-VL[6], despite being a general-purpose multimodal model, underperforms drastically across all metrics. This suggests its current limitations of OCR-free and VLM methods in handling text-centric visual grounding tasks.
Notably, with the provided OCR input context, LLaMA 3.1 achieves competitive performance in predicting the visual grounding indicating that LLMs can effectively reason about spatial relationships when given structured textual context.
We also present a comparative evaluation at the block level in Table 5. Once again, the Region Matching Algorithm maintains superior performance at the block level, emphasizing its consistent advantage in spatially grounded reasoning. The performance gap also reinforces the importance of layout-aware alignment in real-world document QA settings.
Table 5: Block-level grounding evaluation across methods
Method Precision (%) Recall (%) F1-score (%)
Region Matching Algorithm 53.89 68.22 60.21
LLaMA 3.1 (Q+A) 41.72 67.47 51.56
LLaMA 3.1 (Question Only) 32.88 54.46 41.01
4.4 Ablation Study
To understand the impact of supporting multi-block and multi-line regions in the proposed region matching algorithm, we perform a detailed ablation study over the number of allowable blocks and lines during inference that enhances the accuracy.
Effect of Maximum Blocks:
We evaluated performance while varying the maximum number of blocks from 1 to 5 in the region matching algorithm using question + ground truth answer alignment. As shown in Table 6, allowing multiple blocks leads to improved recall at the expense of precision. The optimal F1-score (62.68%) is achieved when allowing up to 2 blocks, indicating that multi-block answers do exist but over-aggregation beyond two blocks introduces noise.
Table 6: Effect of max blocks on region matching performance.
Max Blocks Precision Recall F1-score
1 71.83 61.15 66.06
2 58.67 67.29 62.68
3 55.57 67.66 61.02
4 54.53 68.22 60.61
5 53.89 68.22 60.21
Effect of Maximum Lines:
We further extended the ablation over the number of lines allowed in the answer span. Table 7 and Figure 4 show the variation in precision, recall, and F1-score across iterations. Precision drops steadily as more lines are aggregated, while recall consistently improves, reaching a plateau around 10 lines. The F1-score stabilizes around 69.3%, with the optimal performance achieved at 5 lines.
Table 7: Effect of max lines on region matching performance.
Max Lines Precision Recall F1-score
1 85.68 41.80 56.19
2 75.16 58.20 65.60
3 70.52 66.09 68.23
4 68.28 69.94 69.10
5 66.89 71.96 69.33
6 66.11 72.67 69.24
7 65.79 73.38 69.38
8 65.44 73.58 69.27
9 65.23 73.68 69.20
10 65.06 73.68 69.10
Figure 4: Precision, Recall, and F1-score trends across increasing max line thresholds.
This study highlights a crucial trade-off: precision dominates at lower aggregation levels, while recall improves as the answer spans grow larger. Allowing multi-line, multi-block reasoning significantly enhances performance, but must be carefully bounded to avoid degrading precision. For our task, 2-block and 5-line aggregation offers the most balanced performance.
5 Conclusion and Future works
In this work, we presented DhrishtiKon, a multi-granular visual grounding framework designed to enhance the interpretability and trustworthiness of VQA systems on text-centric images. By integrating multi-lingual OCR, LLMs, and a novel region matching algorithm, we demonstrated significant improvements in grounding accuracy across multiple levels of granularity. Our experiments on a curated dataset derived from CircularsVQA testset showed that mid-level granularities (line-level) yield the best performance for document-based question answering tasks. The ablation studies further validated the effectiveness of multi-block and multi-line reasoning in improving grounding precision and recall. Also, the evaluations highlight that while large vision-language models can produce reasonable semantic predictions, they fall short in precise visual localization tasks without structured grounding mechanisms. Our Region Matching approach demonstrates that traditional alignment-based methods—when combined with robust OCR and textual matching—remain highly competitive for document understanding tasks that demand spatial fidelity.